Suppose you are optimistic on Boeing company. The current market price per share is $50, and you have $5000 of your own money to invest. You decide to borrow an additional $5000 from your broker at an interested rate of 8% per year and invest $10000 in the stock. (a) How far does the price of Boeing company have to fall for you to get a margin call if the maintenance margin is 30%? Assume the price fall happens immediately. (b) Assume the price fall will happen after 1 year. What is your answer to the question above?
Answer #1

a

current price*initial margin-drop in price = maintenance margin*(current price-drop in price)
50*0.5-Drop in price = 0.3*(50-Drop in price)
Drop in price = 14.29
Margin call price = current price-drop in price =50-14.29=35.71

b

After one year initial margin = (old margin-interest rate*amount borrowed)/total amount

=(5000-0.08*5000)/10000

=0.46

current price*initial margin-drop in price = maintenance margin*(current price-drop in price)
50*0.46-Drop in price = 0.3*(50-Drop in price)
Drop in price = 11.43
Margin call price = current price-drop in price =50-11.43=38.57